OpenStudy (anonymous):

What are the x- and y – intercepts of the following equation: y = -0.05x + 16 (16, 0), (0, 320) (0, 16), (32, 0) (0, 16), (320, 0) (16, 0), (-320, 0)

OpenStudy (anonymous):

the method for finding intercepts is, if you think about it you can just substitute. when x = 0 what is the y value, and when y= 0 solve for x

OpenStudy (anonymous):

the simplest one is when x = 0 because that leaves y = 0 +16 so y = 16 now when y = 0 0 = -.05x +16

OpenStudy (baldymcgee6):

and also a function in the form of y=mx+b (like you have) the 'b" is simply your y intercept, (when x=0)

OpenStudy (anonymous):

precisely . which happens when x=0

OpenStudy (anonymous):

do you know how to solve 0=0.05x +16 for x?

OpenStudy (anonymous):

the numbers bellow are the answers i can use i understand what your trien to say but hpw do i get the answer?

OpenStudy (anonymous):

the answers are giving you the ordered pairs for the answer. for example as shown above when x=0 y=16 so the answer for the x intercept is (0,16) and if you solve for when y=0 your answer will be of the form (x value, 0) and that x value is when was solved for when y=0